OpenStudy (anonymous):

Solve using two differemt methods. Explain which method you found to be more efficient. a. 3x-9y=3 6x - 3y= -24

OpenStudy (anonymous):

The two methods are elimination and substitution.

OpenStudy (anonymous):

There is also graphing

OpenStudy (skullpatrol):

What have you tried?

OpenStudy (anonymous):

Nothing yet, i know how to do elimination. But i dont know how to substitution or graphing.

OpenStudy (skullpatrol):

Use elimination and find the answer first. All the methods should give the same answer, right?

OpenStudy (anonymous):

Ok

OpenStudy (skullpatrol):

$$ 3x-9y=3$$ $$ 6x - 3y= -24$$

OpenStudy (anonymous):

I got 3x-6y= -27 but i dont think its rigth is it?

OpenStudy (skullpatrol):

No.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

well, how do i do it?

OpenStudy (anonymous):

Do it simultaneously.

OpenStudy (skullpatrol):

You said you know how to use elimination. Try multiplying the first equation by $$-2$$ and add it to the second equation...

OpenStudy (skullpatrol):

$$ (3x-9y=3) (-2)$$ 6x - 3y= -24

OpenStudy (anonymous):

\[3x-9y=3 [1]\] \[6x - 3y= -24 [2]\] \[2\times [1]\] \[6x-18y=6 [3]\] \[[2]-[3]\] \[(6x - 3y) - (6x-18y)= (-24)-(6)\]

OpenStudy (anonymous):

\[-3y+18y=-30\] The x's cancel out. Now you can solve for y. Once you solve for y, you can then substitute the y component that you just found, into any one of the equations above to find your x-value.

OpenStudy (anonymous):

If you need further assistance, I will be happy to elaborate it in more simple terms.